# 2cm2cm1cm1cm

### v1

@@ 2.4 Fiber Decomposition of the Hamiltonian
By explicit computations, exactly like in [22, 21], we show that the conjugation of the Hamiltonian $H\in\mathcal{B}(\mathfrak{H})$ with the unitary transformation $U$ of Equation (24) is a decomposable operator on the direct integral $L^{2}(\mathbb{T}^{2},\mathcal{H})$. To state this result precisely, we need preliminary definitions allowing to define the so-called “fiber Hamiltonians”, or “fibers” for short, $A(k)\in\mathcal{B}(\mathcal{H})$ of $UHU^{\ast}$ at total quasi-momenta $k\in\mathbb{T}^{2}$. In fact, the mapping $k\mapsto A(k)$ defines an element of the von Neumann algebra
@@
of (equivalence classes of) strongly measurable functions $\mathbb{T}^{2}\rightarrow\mathcal{B}(\mathcal{H})$. Recall that the (unique) norm of this von Neumann algebra (as a $C^{\ast}$-algebra) is the essential supremum (with respect to the measure $\nu$ on the torus):
$\left\|B\right\|_{\infty}\doteq\underset{k\in\mathbb{T}^{2}}{\mathrm{ess-sup}}\,\left\|B\left(k\right)\right\|_{\mathrm{op}}<\infty\text{ },$
where $\left\|\cdot\right\|_{\mathrm{op}}$ stands for the operator norm. In fact, $A(\cdot)$ is even a continuous function on the torus $\mathbb{T}^{2}$ as compact metric space. See Section 5.3 for more details.
@@
$\mathfrak{d}\left(k\right)\left(p\right)$ $\doteq$ $\mathfrak{\hat{p}}\left(k+p\right)\ ,$ (29)
@@
Recall that (the $(2\pi,2\pi)$-periodic function) $\mathfrak{\hat{p}}$ is the Fourier transform of $\mathfrak{p}$, which is the function defining the operator $c_{y}^{\ast}$ in (11), representing the creation of fermion pairs in the model.
@@
where $M_{\mathfrak{f}(k)}$ stands for the multiplication operator by $\mathfrak{f}(k)\in C(\mathbb{T}^{2})$ and $P_{x}$ is the orthogonal projection onto the one-dimensional subspace $\mathbb{C}\mathfrak{\hat{e}}_{x}\subseteq L^{2}(\mathbb{T}^{2})$. Note that the infinite sum defining the bounded operator $B_{1,1}(k)$ is absolutely convergent, for the function $\mathrm{u}:\mathbb{Z}^{2}\rightarrow\mathbb{R}$ is, by assumption, absolutely summable. See (7). We define next
@@
$A\left(\cdot\right)\equiv A\left(\mathrm{U},\cdot\right)\doteq\begin{pmatrix}A_{1,1}\left(\mathrm{U},\cdot\right)&A_{1,2}\left(\cdot\right)\\[5.0pt] A_{2,1}\left(\cdot\right)&A_{2,2}\left(\cdot\right)\end{pmatrix}\in L^{\infty}\left(\mathbb{T}^{2},\mathcal{H}\right)\ ,$ (40)
which is meanwhile the fiber decomposition of the operator $UHU^{\ast}$:
@@
The conjugation of $H$ by $U$ (24) is decomposable and has $A(\cdot)$ as its fibers, that is,
$UHU^{\ast}={\int_{\mathbb{T}^{2}}^{\oplus}}A(k)\,\nu(\mathrm{d}k)\ .$
Proof. This is proven from explicit computations which are almost the same as those done in [22, 21]. We omit the details.
@@ 3 Main Results
$\mathrm{u}:\mathbb{Z}^{2}\rightarrow\mathbb{R}_{0}^{+}\ ,\quad\mathfrak{p}:\mathbb{Z}^{2}\rightarrow\mathbb{R}\ ,\quad\text{and}\quad\upsilon:\mathbb{Z}^{2}\rightarrow\mathbb{R}$
that are absolutely summable and invariant with respect to $90^{\circ}$-rotations. Observe additionally that the function $\mathfrak{p}$ is required to be exponentially decaying, that is, $\mathrm{e}^{\alpha_{0}|\cdot|}\mathfrak{p}$ is absolutely summable for some $\alpha_{0}>0$. See Equations (7), (12) and (14). All details of the Hamiltonian, like the precise choice of its parameters and functions, are not explicitly mentioned in our discussions or statements below, unless it is important for clearness. There is however one important condition to clarify:
While some of our results can be obtained without any other restriction, frequently we fix the parameter $h_{b}$ in the interval $[0,1/2]$. This choice physically means that the boson is heavier than two fermions. As already discussed above, the assumption is perfectly justified when one views the two fermions and the boson of the model as being electrons or holes and a bipolaron, respectively, in a cuprate. In fact, polarons (and thus bipolarons) are charge carriers that are self-trapped inside a strong and local lattice deformation that surrounds them, caused by electrostatic interactions between the carriers and the lattice. A priori, such (strong and local) lattice deformations can barely move, that is, their effective mass is huge. See, e.g., [28, 29, 30]. This is coherent with the assumption of a large mass of JT bipolarons in copper oxides [31], similar to JT polarons [33].
@@ 3.1 Spectral Properties
Proof. The theorem is a combination of Theorems 4.7, 4.8, 4.17 and 4.19 together with Propositions 4.1, 4.18 and Corollary 4.5 (see (118)).
@@
If $\mathfrak{p}\in\mathbb{C}\mathfrak{e}_{0}$, i.e., $r_{\mathfrak{p}}=0$, then Theorem 3.1 (iii) remains true, but not uniformly in $\mathrm{U}\in\mathbb{R}_{0}^{+}$. That is, in this case, one only has:
@@
As is explained in [22, 21], the eigenvalue $\mathrm{E}\left(\mathrm{U},k\right)$ given by Theorem 3.1 is associated with the formation of dressed bound fermion pairs with total quasi-momentum $k\in\mathbb{T}^{2}$. These pairs are generally exponentially localized, thanks to Theorem 3.1 (iii), which basically implies that the two fermions move together confined within some small ball, that is, they are tightly bound in space, provided $\hat{\upsilon}(0)\neq 0$. When $\mathfrak{p}(z)\neq 0$ for some $z\neq 0,$ or, equivalently, $r_{\mathfrak{p}}>0$, the size of the small does not depend upon the Hubbard coupling constant $\mathrm{U}$ and a very large $\mathrm{U}\gg 1$ only prevents two fermions from occupying the same lattice site. The condition $\mathfrak{p}(z)\neq 0$ for some $z\neq 0$ or, equivalently, $r_{\mathfrak{p}}>0$, is therefore pivotal to get (Cooper) fermion pairs, the natural candidates for superconducting charge carriers, in presence of strong on-site Coulomb repulsions, like in cuprates [23, 24].
For the usual (i.e., non-extended) Hubbard interaction ($\mathrm{u}=0$) and one-range [^6] It means here that $c_{y}^{\ast}\doteq\sum_{z\in\mathbb{Z}^{2}:\left|z\right|\leq 1}\,\mathfrak{p}(z)(a_{y+z,\uparrow}^{\ast}\,a_{y,\downarrow}^{\ast}+a_{y+z,\uparrow}^{\ast}\,a_{y-z,\downarrow}^{\ast})$. creation operators $c_{y}^{\ast}$ of fermion pairs (in this case (12) holds true for all $\alpha_{0}\in\mathbb{R}_{0}^{+}$), note that a weak form of pair localization was previously shown in the ground state. See for instance [22, Theorem 3 and Proposition 13]. In this particular case, estimates of $\mathrm{E}\left(\mathrm{U},k\right)$ and $\Psi(\mathrm{U},k)$ are known for large $\mathrm{U}\gg 1$. See, e.g., [22, Theorem 4, Corollary 5, Theorem 16]. Recall that the aim in [22, 21] was to show the existence of $d$- and $p$-wave pairings in the ground state for some physically sound model, and not the systematic study of a general class of models. In [21] we conjecture that such dressed bound fermion pairs represent the charge carriers below the pseudogap temperature in cuprates.
@@
$\left\{\mathrm{E}\left(\mathrm{U},k\right):k\in\mathbb{T}^{2}\right\}\cup\left(0,8\epsilon\right)\subseteq\sigma\left(H\right)\ ,$
@@ 3.2 Dispersion Relation of Dressed Bound Fermion Pairs
with $\mathsf{GL}_{2}\left(\mathbb{R}\right)\subseteq\mathsf{M}_{2}\left(\mathbb{R}\right)$ being the set of invertible $2\times 2$ matrices with real coefficients. As $\mathsf{GL}_{2}(\mathbb{R})\subseteq\mathsf{M}_{2}(\mathbb{R})$ is an open set (see [44, Theorem 1.4]), it then follows that
@@
of free bosons (bipolarons for cuprates). By turning on the fermion-boson-exchange interaction, the dispersion relation of dressed bound fermion pairs with lowest energy can strongly deviate from $\mathfrak{b}$, the unperturbed one. Recall for instance that $\mathfrak{b}$ describes bosons with a very large mass as compared to the effective mass of electrons or holes in cuprates. However, as shown in [21], for typical parameters of the cuprate LaSr 214, the effective mass of the bound pair (with dispersion relation $\mathrm{E}$) is comparable to the mass of electrons or holes. This solves the so-called “large mass paradox” of the microscopic theory of cuprate superconductors, based on some kind mechanism involving bipolarons. For more details, see [21] and references therein.
@@ 3.3 Quantum Scattering
Scattering in quantum mechanics constitutes a well-established mathematical theory aiming at analyzing the behavior of quantum systems at large times. To this end, a reference (or free) Hamiltonian $Y$ is chosen and the dynamics $(\mathrm{e}^{itX})_{t\in\mathbb{R}}$ of the quantum system driven by the (full) Hamiltonian $X$ is compared at large (negative and positive) times to $(\mathrm{e}^{itY})_{t\in\mathbb{R}}$. In fact, scattering theory can be viewed as a kind of perturbation theory for the absolutely continuous spectrum of $X$. See, e.g., [34, Chapter X]. For standard textbooks explaining in detail the scattering theory, we recommend [35, 36, 37]. Below, for the reader’s convenience, we shortly recall definitions that are relevant here.
@@
when it exists. See for instance [37, Definition 1.3]. When $\mathcal{Y}=\mathcal{X}$ and $J=\mathfrak{1}$, like in [37, Definition 1.1], we use the shorter notation
@@
In case the above wave operators exist, they are partial isometries [35, Proposition 1, Sect. XI.3]. They are said to be complete when
@@
See [35, p. 19, Sect. XI.3]. Similarly, in the general case, $W^{\pm}\left(X,Y;J\right)$ are said to be complete whenever
@@
See [35, p. 35, Sect. XI.3]. The corresponding scattering operator is equal to
@@
It leads to the scattering matrix (or simply $S$-matrix) in a representation where $Y$ is diagonal, because the scattering operator commutes with $Y$. See [37, Equation (1.12)].
@@
$UHU^{\ast}={\int_{\mathbb{T}^{2}}^{\oplus}}A(k)\,\nu(\mathrm{d}k)$

### v2

@@ 2.4 Fiber Decomposition of the Hamiltonian
By explicit computations, exactly like in [22, 21], we show that the conjugation of the Hamiltonian $H\in\mathcal{B}(\mathfrak{H})$ with the unitary transformation $\mathbb{U}$ of Equation (26) is a decomposable operator on the direct integral $L^{2}(\mathbb{T}^{2},\mathcal{H})$. To state this result precisely, we need preliminary definitions allowing to define the so-called “fiber Hamiltonians”, or “fibers” for short, $A(k)\in\mathcal{B}(\mathcal{H})$ of $\mathbb{U}H\mathbb{U}^{\ast}$ at total quasi-momenta $k\in\mathbb{T}^{2}$. In fact, the mapping $k\mapsto A(k)$ defines an element of the von Neumann algebra [^10] The (unique) norm of this $C^{\ast}$-algebra is the essential supremum with respect to the measure $\nu$ on the torus, see (164).
@@
of (equivalence classes of) strongly measurable functions $\mathbb{T}^{2}\rightarrow\mathcal{B}(\mathcal{H})$. See Section 5.4 for more details.
@@
$\mathfrak{d}\left(k\right)\left(p\right)$ $\doteq$ $\mathfrak{\hat{p}}_{1}\left(k+p\right)+\mathfrak{\hat{p}}_{2}\left(k/2+p\right)\ ,$ (34)
@@
Recall that (the $(2\pi,2\pi)$-periodic function) $\mathfrak{\hat{p}}_{1}$ and $\mathfrak{\hat{p}}_{2}$ are the Fourier transform of $\mathfrak{p}_{1}$ and $\mathfrak{p}_{2}$, which are the functions defining the operator $c_{y}^{\ast}$ in (11), representing the creation of fermion pairs in the model. Note also from (13) that $\mathfrak{d}(k)\neq 0$ for all $k\in\mathbb{T}^{2}$. Indeed, using $\mathfrak{p}_{2}(z)\doteq 0$ for $z\notin 2\mathbb{Z}$ as well as (31) and (34),
$\mathfrak{d}\left(k\right)=\mathcal{F}\left[\mathrm{e}^{ik\cdot x}\mathfrak{p}_{1}\left(x\right)+\mathrm{e}^{i\frac{k}{2}\cdot x}\mathfrak{p}_{2}\left(x\right)\right]\ ,$ (36)
where $\mathrm{e}^{ik\cdot x}\mathfrak{p}_{\sharp}(x)$ stands for the function $x\mapsto\mathrm{e}^{ik\cdot x}\mathfrak{p}_{\sharp}(x)$ with $\sharp\in\{1,2\}$.
@@
where $M_{\mathfrak{f}(k)}$ stands for the multiplication operator by $\mathfrak{f}(k)\in C(\mathbb{T}^{2})$ and $P_{x}$ is the orthogonal projection onto the one-dimensional subspace $\mathbb{C}\mathfrak{\hat{e}}_{x}\subseteq L^{2}(\mathbb{T}^{2})$. Note that the infinite sum defining the bounded operator $B_{1,1}(k)$ is absolutely convergent, for the function $\mathrm{u}:\mathbb{Z}^{2}\rightarrow\mathbb{R}$ is, by assumption, absolutely summable. See (7).
We define next
@@
$A\left(\cdot\right)\equiv A\left(\mathrm{U},\cdot\right)\doteq\begin{pmatrix}A_{1,1}\left(\mathrm{U},\cdot\right)&A_{1,2}\left(\cdot\right)\\[5.0pt] A_{2,1}\left(\cdot\right)&A_{2,2}\left(\cdot\right)\end{pmatrix}\in L^{\infty}\left(\mathbb{T}^{2},\mathcal{B}(\mathcal{H})\right)$ (46)
(see Lemma 4.1), which is meanwhile the fiber decomposition of the operator $\mathbb{U}H\mathbb{U}^{\ast}$:
@@
The conjugation of $H$ by $\mathbb{U}$ (26) is decomposable and has $A(\cdot)$ as its fibers, that is,
$\mathbb{U}H\mathbb{U}^{\ast}={\int_{\mathbb{T}^{2}}^{\oplus}}A(k)\,\nu(\mathrm{d}k)\ .$
Proof. This is proven from explicit computations which are almost the same as those done in [22, 21]. We postpone the details of this calculation to Section 4.2.
@@ 3 Main Results
$\mathrm{u}:\mathbb{Z}^{2}\rightarrow\mathbb{R}_{0}^{+}\ ,\quad\mathfrak{p}_{1}:\mathbb{Z}^{2}\rightarrow\mathbb{R}\ \ ,\quad\mathfrak{p}_{2}:\mathbb{Z}^{2}\rightarrow\mathbb{R}\quad\text{and}\quad\upsilon:\mathbb{Z}^{2}\rightarrow\mathbb{R}$
(with $\mathfrak{p}_{2}(z)\doteq 0$ for $z\notin 2\mathbb{Z}$) that are absolutely summable and invariant with respect to $90^{\circ}$-rotations. Observe additionally that the functions $\mathfrak{p}_{1}$ and $\mathfrak{p}_{2}$ are required to be exponentially decaying, that is, $\mathrm{e}^{\alpha_{0}|\cdot|}\mathfrak{p}_{1}$ and $\mathrm{e}^{\alpha_{0}|\cdot|}\mathfrak{p}_{2}$ are absolutely summable for some $\alpha_{0}>0$. See Equations (7), (12) and (16). All details of the Hamiltonian, like the precise choice of its parameters and functions, are not explicitly mentioned in our discussions or statements below, unless it is important for clearness. There is however one important condition to clarify:
While some of our results can be obtained without any other restriction, frequently we fix the parameter $h_{b}$ in the interval $[0,1/2]$. This choice physically means that the boson is heavier than two fermions. As already discussed above, the assumption is perfectly justified when one views the two fermions and the boson of the model as being electrons or holes and a bipolaron, respectively, in a cuprate. In fact, polarons (and thus bipolarons) are charge carriers that are self-trapped inside a strong and local lattice deformation that surrounds them, caused by electrostatic interactions between the carriers and the lattice. A priori, such (strong and local) lattice deformations can barely move, that is, their effective mass is huge. See, e.g., [34, 35, 36]. This is coherent with the assumption of a large mass of JT bipolarons in copper oxides [37], similar to JT polarons [39]. See also Section 5.1 for more details.
@@ 3.1 Spectral Properties
Proof. The theorem is a combination of Theorems 4.8, 4.9, 4.18 and 4.20 together with Propositions 4.2, 4.19 and Corollary 4.6 (see (134)).
@@
If $\mathfrak{p}_{1}=\mathfrak{p}_{2}\in\mathbb{C}\mathfrak{e}_{0}$,i.e., $r_{\mathfrak{p}}=0$, then Theorem 3.1 (iii) remains true, but not uniformly in $\mathrm{U}\in\mathbb{R}_{0}^{+}$. That is, in this case, one only has:
@@
As is explained in [22, 21], the eigenvalue $\mathrm{E}\left(\mathrm{U},k\right)$ given by Theorem 3.1 is associated with the formation of dressed bound fermion pairs with total quasi-momentum $k\in\mathbb{T}^{2}$. These pairs are generally exponentially localized, thanks to Theorem 3.1 (iii), which basically implies that the two fermions move together confined within some small ball, that is, they are tightly bound in space, provided $\hat{\upsilon}(0)\neq 0$. When $\mathfrak{p}_{1}(z)\neq 0$ or $\mathfrak{p}_{2}(z)\neq 0$ for some $z\neq 0,$ or, equivalently, $r_{\mathfrak{p}}>0$, the size of the small does not depend upon the Hubbard coupling constant $\mathrm{U}$ and a very large $\mathrm{U}\gg 1$ only prevents two fermions from occupying the same lattice site. The condition $\mathfrak{p}_{1}(z)\neq 0$ or $\mathfrak{p}_{2}(z)\neq 0$ for some $z\neq 0$ or, equivalently, $r_{\mathfrak{p}}>0$, is therefore pivotal to get (Cooper) fermion pairs, the natural candidates for superconducting charge carriers, in presence of strong on-site Coulomb repulsions, like in cuprates [29, 30].
For the usual (i.e., non-extended) Hubbard interaction ($\mathrm{u}=0$) and one-range [^12] It means here that $c_{y}^{\ast}\doteq{\sum_{\left|z\right|\leq 1}}\,\big{(}\mathfrak{p}_{1}\left(z\right)a_{y+z,\uparrow}^{\ast}\,a_{y,\downarrow}^{\ast}+\mathfrak{p}_{2}\left(2z\right)a_{y+z,\uparrow}^{\ast}\,a_{y-z,\downarrow}^{\ast}\big{)}$. creation operators $c_{y}^{\ast}$ of fermion pairs (in this case (12) holds true for all $\alpha_{0}\in\mathbb{R}_{0}^{+}$), note that a weak form of pair localization was previously shown in the ground state. See for instance [22, Theorem 3 and Proposition 13]. In this particular case, estimates of $\mathrm{E}\left(\mathrm{U},k\right)$ and $\Psi(\mathrm{U},k)$ are known for large $\mathrm{U}\gg 1$. See, e.g., [22, Theorem 4, Corollary 5, Theorem 16]. Recall that the aim in [22, 21] was to show the existence of $d$- and $p$-wave pairings in the ground state for some physically sound model, and not the systematic study of a general class of models. In [21] we conjecture that such dressed bound fermion pairs represent the charge carriers below the pseudogap temperature in cuprates.
@@
$\sigma\left(H\right)\cap\left(-\infty,8\epsilon\right]=\left\{\mathrm{E}\left(\mathrm{U},k\right):k\in\mathbb{T}^{2}\right\}\cup\left(0,8\epsilon\right)\ ,$
@@ 3.2 Dispersion Relation of Dressed Bound Fermion Pairs
with $\mathsf{GL}_{2}\left(\mathbb{R}\right)\subseteq\mathsf{M}_{2}\left(\mathbb{R}\right)$ being the set of invertible $2\times 2$ matrices with real coefficients. As $\mathsf{GL}_{2}(\mathbb{R})\subseteq\mathsf{M}_{2}(\mathbb{R})$ is an open set (see [50, Theorem 1.4]), it then follows that
@@
of free bosons (bipolarons for cuprates).
By turning on the fermion-boson-exchange interaction, the dispersion relation of dressed bound fermion pairs with lowest energy can strongly deviate from $\mathfrak{b}$, the unperturbed one. Recall for instance that $\mathfrak{b}$ describes bosons with a very large mass as compared to the effective mass of electrons or holes in cuprates. However, as shown in [21], for typical parameters of the cuprate LaSr 214, the effective mass of the bound pair (with dispersion relation $\mathrm{E}$) is comparable to the mass of electrons or holes. This is a consequence of the mass of charge carriers calculated in [52], and the fact that a large effective mass of dressed bound fermion pairs and a high fermion-pair depletion [^13] I.e., the fermionic component of the dressed bound fermion pairs is very small in comparison with the bipolaronic component., close to $90\%$ as measured in [53], is not compatible with our model. This solves the so-called “large mass paradox” of the microscopic theory of cuprate superconductors, based on some kind mechanism involving bipolarons. For more details, see [21] and references therein.
In fact, the effective mass of dressed bound fermion pairs, or more generally its (effective) mass tensor, depends strongly on the coupling function $\hat{\upsilon}$ near its maximum. Bearing in mind Definitions 3.7 and 3.8, one can therefore provide via Theorems 3.6 and 3.8 not only qualitative but also quantitative information, which is important for describing the physical behaviour of fermionic pairs formed in this way by means of a bosonic field. A natural question is then to study its scattering properties and this is precisely what we propose to do in the next section.
@@ 3.3 Quantum Scattering
Scattering in quantum mechanics constitutes a well-established mathematical theory aiming at analyzing the behavior of quantum systems at large times. To this end, a reference (or free) Hamiltonian $Y$ is chosen and the dynamics $(\mathrm{e}^{itX})_{t\in\mathbb{R}}$ of the quantum system driven by the (full) Hamiltonian $X$ is compared at large (negative and positive) times to $(\mathrm{e}^{itY})_{t\in\mathbb{R}}$. In fact, scattering theory can be viewed as a kind of perturbation theory for the absolutely continuous spectrum of $X$. See, e.g., [40, Chapter X]. For standard textbooks explaining in detail the scattering theory, we recommend [41, 42, 43]. Below, for the reader’s convenience, we shortly recall definitions that are relevant here.
@@
when it exists. See for instance [43, Definition 1.3]. When $\mathcal{Y}=\mathcal{X}$ and $J=\mathfrak{1}$, like in [43, Definition 1.1], we use the shorter notation
@@
In case the above wave operators exist, they are partial isometries [41, Proposition 1, Sect. XI.3]. They are said to be complete when
@@
See [41, p. 19, Sect. XI.3].
Similarly, in the general case, $W^{\pm}\left(X,Y;J\right)$ are said to be complete whenever
@@
See [41, p. 35, Sect. XI.3]. The corresponding scattering operator is equal to
@@
It leads to the scattering matrix (or simply $S$-matrix) in a representation where $Y$ is diagonal, because the scattering operator commutes with $Y$. See [43, Equation (1.12)].
@@
$\mathbb{U}H\mathbb{U}^{\ast}={\int_{\mathbb{T}^{2}}^{\oplus}}A(k)\,\nu(\mathrm{d}k)$